# Measuring the Hidden Cost of Data Valuation through Collective Disclosure

### v1

@@ 1 Introduction
The idea of data dividends (Feygin et al., 2021)—distributing a share of value generated from data back to contributors—has drawn attention as a mechanism to make data economies more inclusive. A way to ground such dividends is through data valuation, in which each point’s contribution to predictive performance is quantified. Recent methods such as data Shapley value (DSV) (Jia et al., 2019b), which originated from cooperative game theory, have been applied to data summarization and efficient acquisition (Ghorbani and Zou, 2019). These methods show that only a fraction of the data is needed for high utility, but nonetheless require access to the entire dataset to compute the marginal contributions—meaning that each individual must first contribute their data, even if their eventual payout is negligible. We refer to this as the hidden cost of data valuation.
Following calls for data trusts and unions as institutional forms of collective governance (Delacroix and Lawrence, 2019), we frame our approach as a form of algorithmic collective action in which contributors act collectively through a Data Union (DU) that sets group data disclosure policies, shifting how value and costs are distributed. While recent work shows that differentially-private mechanisms  Dwork (2006) can diminish the leverage of collective action in gradient-based training (Solanki et al., 2025), we instead use DP positively as a tool for disclosure and value control.
@@
Figure 1: Illustration of the Information Disclosure Game. A Data Union (DU) holds a private dataset and releases information to a Data Consumer (DC) by adding Laplacian noise to data points under $\epsilon$-differential privacy. The DC incrementally acquires these noisy version of the data points, denoises them using an average and train a model to reach a utility target. In this work, we focus on non-parametric $k$-Nearest Neighbors (kNN) and SBERT embeddings Reimers and Gurevych (2019).
We formalize this interaction as an Information Disclosure Game (IDG), a form of Stackelberg game Simaan and Cruz Jr (1973); Fudenberg and Tirole (1991) in which the DU acts as leader by choosing disclosure rules, and the DC acts as follower by optimizing acquisitions given those rules (Figure 1). This framework allows to make explicit the acquisition cost that valuation imposes on the DC and how collective disclosure policies shape their distribution across members. We instantiate the model by a $k$-Nearest Neighbors (kNN), in which the relation between individual points and predictive utility is direct. Simulating the DC strategies guided by Shapley values and by multi-armed bandit exploration, we test it on a Yelp review helpfulness prediction task using text embeddings.
@@
First in Section 2, we review the related work on data valuation and differential privacy before detailing in Section 3 the IDG framework and the objectives of DU and DC. Afterwards in Section 4, we present our empirical evaluation followed by a discussion and conclusion.
@@ 2 Background
In cooperative game theory, an imputation is a way to distribute the total value of a game among its players Osborne and Rubinstein (1994). In the context of data valuation, each data point in a dataset $D$ is treated as a player and a value function $v:2^{D}\to\mathbb{R}$ is used to the utility (e.g., accuracy or loss reduction) of any subset $D^{\prime}\subseteq D$. The objective is to find an imputation that fairly allocates the total value $v(D)$ among individual data points based on their contributions. Several approaches have been explored for data valuation—each with different fairness, computational and interpretability trade-offs—including Data Shapley Ghorbani and Zou (2019); Jia et al. (2019b), Data Banzhaf Wang and Jia , Beta Shapley Kwon and Zou and the Core Yan and Procaccia .
@@
in which $D^{\prime}$ is a subset of $D\setminus\{z_{i}\}$ sampled uniformly at random and $v(D^{\prime})$ is the value function evaluated on $D^{\prime}$. The Shapley value thus quantifies the expected incremental contribution of each data point across all possible subsets. The uniqueness of the Shapley value can be derived from satisfying four foundational axioms: Efficiency (i.e., the total value is distributed among all players), Symmetry (i.e., identical contributions are rewarded equally), Dummy (i.e., players that contribute nothing receive zero value) and Linearity (i.e., the Shapley values from two games can be combined linearly). One application is data pricing in marketplaces Jia et al. ; Pei ; Tian et al. ; Xia et al. , in which Shapley-based frameworks assign higher prices to points that contribute across many subsets. Shapley values can also inform the data selection task, by identifying examples most useful for training smaller high-quality sets. Their effectiveness depends on the structure of the utility function and faces the challenge of high computational cost, with approximations such as G-Shapley Ghorbani and Zou (2019); Yoon et al. (2020) proposed to improve scalability. However, for $k$-NN classifiers, an exact Shapley formulation exists Jia et al. (2019a) with complexity $O(N\log N)$, making valuation tractable in this setting.
@@
Differential Privacy (DP) Dwork (2006) provides a mathematically rigorous privacy framework by ensuring that the inclusion or exclusion of any individual in a dataset does not significantly alter the probability distribution of outputs. For instance, the trade-off between privacy and utility can be controlled through noise addition, thus reducing the re-identification risk and providing provable privacy guarantees. Formally, let $D=\{z_{1},z_{2},\dots,z_{n}\}$ be a dataset consisting of $n$ individual data points. A randomized mechanism $M$ satisfies $(\epsilon,\delta)$-differential privacy if, for any dataset $D$, any data point $z_{i}\in D$, and any subset of possible outputs $O$, the following holds:
$\Pr[M(D)\in O]\leq e^{\epsilon}\Pr[M(D\setminus\{z_{i}\})\in O]+\delta,$
in which $\epsilon$ represents the privacy loss and $\delta$ allows for a small probability that strict $\epsilon$-differential privacy does not hold. Beyond protection, privacy also concerns the incentives individuals face when deciding whether to share their data truthfully. From a mechanism design perspective, the Revelation Principle Myerson (1983) states that outcomes achievable through indirect strategies can also be implemented by truthful revelation. Differential privacy makes such mechanisms approximately truthful: in McSherry and Talwar (2007), an $\varepsilon$-DP mechanism bounds each agent’s influence on the outcome, so the gain from lying is at most $O(\varepsilon)$.
@@ 3 Information Disclosure Game
We model the interaction between two agents: the DU and the DC. As a member-run collective, the DU manages pooled data with the mission of ensuring fair valuation while accounting for concerns such as privacy. Meanwhile, the DC seeks to optimize utility—often by acquiring data points at minimal cost. These interactions are formalized as a two-phase Stackelberg game, in which the DU acts as a leader by setting how to disclose information and at what cost, and the DC, as a follower, responds through strategic data acquisition.
@@
This approach provides a straightforward pricing mechanism but has significant limitations. Indeed, because disclosure of a data point is binary—either withheld or complete—many contributors receive no compensation. Incentivizing broader acquisition requires lowering the price of high-value points, which penalizes those who contribute the most. Moreover, releasing data at fixed prices locks in value at a single point in time and weakens privacy guarantees, limiting the DU’s ability to balance inclusiveness and protection.
@@
To address these challenges, we depart from per-point pricing and instead model an iterative disclosure process under DP. At each round $t$, the DC selects a subset $S_{t}$ of data points to query. Each query consumes a fixed privacy budget $\epsilon$ and returns a noisy version of the data. The DU sets $(\epsilon,T_{\max})$, the per-query budget and the per-point query cap, which together determine a maximum per-point spend $B_{\max}=T_{\max}\epsilon$. The DC’s cumulative spend is $B=\epsilon\sum_{t=1}^{T}|S_{t}|,\quad T\leq T_{\max}$.
The DC’s objective is to minimize its cumulative budget spent:
@@
$\max_{(\epsilon,T_{\max})}\;\min_{\{S_{t}\}}\epsilon\sum_{t=1}^{T}|S_{t}|$ (3)
with the goal of spreading spend more evenly across members while managing privacy–utility trade-offs. Compared to the pricing game in Equation 1, partial disclosure replaces per-point ownership costs with iterative access charges. This formulation makes the hidden acquisition cost of valuation explicit in the DC’s budget and allows the DU to shape acquisition behaviors without penalizing the most valuable contributors.
@@ 4 Experiments on Review Helpfulness Prediction
Review helpfulness prediction provides a natural testbed for data valuation as platforms such as Yelp collect reviews not only for sentiment analysis but also to assess their quality. Indeed, unhelpful reviews may ultimately be discarded, yet they are still necessary to identify and reward helpful ones. This creates a setting in which individual contributions to predictive performance are both explicit and uneven. We use the Yelp dataset Asghar (2016) following the setup of Bilal and Almazroi (2023), who report a $k$-NN accuracy of 59.6%. Using pretrained text embeddings and $k$-NN, our best configuration achieves 66.0% test accuracy and 65.2% F1-score, bringing non-parametric performance closer to fine-tuned transformer models while preserving point-level interpretability needed for valuation. As a sanity check, we replicated the acquisition experiment of Jia et al. (2019b) and verified that Shapley-based selection outperforms random sampling. The corresponding curve, along with additional details on the experimental setup, is reported in Appendix D.
@@
in which $\Delta_{j}=b_{j}-a_{j}$ denotes the global sensitivity of feature $j$ after bounding and $\epsilon_{j}$ is the allocated budget for that feature. This design is similar in spirit to a per-feature local DP mechanism Cormode et al. (2018); Wang et al. (2019), in that each coordinate is privatized independently, but here it is applied in a centralized setting by the DU. We assume the sensitivity interval $a_{j},b_{j}]$ is fixed in advance and does not vary significantly across releases, so the noise scale is determined by $\Delta_{j}/\epsilon_{j}$. On the DC side, each noisy version of a point is averaged to form a denoised estimate. After observing $t$ noisy versions of point $i$, its center is computed incrementally as $\hat{x}_{i}^{(t)}=\frac{1}{t}\sum_{k=1}^{t}\tilde{x}_{i}^{(k)}$. The averaging strategy steadily improves fidelity as noisy samples accumulate, with error decreasing and correlation increasing non-linearly across iterations (see Appendix B for details).
@@ 4.1 Data Selection Using Random and Shapley-based Strategies
In this baseline, the DC commits to a random subset of data points for all iterations with the results displayed in Figure 3 being averaged over 10 random permutations. As shown in Figure 3, this approach fails to reach the target utility within 100 iterations unless nearly 100% of the dataset is acquired, which confirms that a purely random acquisition strategy from the DU is not viable under budget constraints.
@@ 4.2 Data Selection using $n$-Armed Bandits
The utility of a queried point is defined as the fraction of neighbors in $k$-NN that share its label. See Algorithm 1 in Appendix for implementation details.
@@
Identifying an equilibrium can be seen as a hyperparameter search, since the interaction between the DU’s release policy and the DC’s acquisition strategy does not admit a closed-form solution. To realize this, we conducted a grid search over two key parameters: the maximum budget per point (set by the DU) and the exploration coefficient $c$ (in UCB). These jointly determine budget usage and the diversity of selected points. Results show that the bandit reliably reaches the utility threshold except when the per-point budget is too small or exploration is disabled; below a budget of 20, success is rare and often due to lucky point selection. Gini analysis of budget allocations confirms that exploration promotes more balanced spending across points and increases the likelihood of success. The correlation between Q values and Shapley values grows with budget but remains moderate overall (Appendix E).Finally, evaluations based on denoised centers yielded test accuracy comparable to or exceeding that obtained from the original data. For instance, a configuration with a per-point budget of 50 reached 69.9% surpassing the 66% benchmark.
@@ 5 Discussion & Conclusion
We proposed a game-theoretic framework for data valuation that integrates privacy and inclusiveness concerns through an information disclosure game between a Data Union (DU) and a data consumer (DC). Rather than assuming direct access to data, our approach models acquisition as an iterative process using a differential private mechanism to enforce a cost on data valuation. On Yelp review helpfulness prediction, Shapley-based strategies remained effective under noise but required a minimal budget—about 10 bootstrap iterations—to reliably identify high-value points. This demonstrates that Shapley valuation entails an inherent cost of exploration in our setting. Similarly, a multi-armed bandit strategy reached the utility target with comparable budgets, reinforcing that valuation itself imposes exploration costs. Gini analysis further showed that as budgets grow, the DC allocates resources more evenly across contributors, increasing inclusiveness.
While our experiments focused on $k$-NN for computation efficiency, an important direction for future work is extending the framework to differentiable models. Gradient-based methods such as G-Shapley Ghorbani and Zou (2019) already show that Shapley values can be approximated from changes in gradients, suggesting ways to scale valuation beyond non-parametric models. In parallel, approaches like DP-SGD Abadi et al. (2016) demonstrate that DP can be enforced by adding noise to gradients rather than directly to points. Bridging these lines of work would provide a better privacy-utility trade-off for both the DU and DC. Nonetheless, our results suggest that in a DU setting, reaching target utility requires a minimum level of budget spread. Thus, our framework implies that a minimum dividend should be guaranteed to all members, regardless of individual Shapley value. In the context of reviews, this means that every contributor would receive at least some share of value, even if their individual review is ultimately deemed unhelpful, aligning the incentives of all members with the collective outcome. Otherwise, excluding some contributors increases the incentive to form data unions that adopt adversarial disclosure strategies (e.g., prioritizing low-value reviews or injecting excess noise), which may in turn hinder the data consumer’s ability to reach its utility target.

### v2

@@ 1 Introduction
The idea of data dividends (feygin2021data)—distributing a share of the value generated from data back to contributors—has drawn attention as a mechanism to make data economies more inclusive. A way to ground such dividends is through data valuation, in which each point’s contribution to predictive performance is quantified. Recent methods such as data Shapley value (DSV) (jia2019towards), which originated from cooperative game theory, have been applied to data summarization and efficient acquisition (ghorbani2019data). These methods show that only a fraction of the data is needed for high utility, but nonetheless require access to the entire dataset to compute the marginal contributions—meaning that each individual must first contribute their data, even if their eventual payout is negligible. We coin this as the hidden cost of data valuation.
Following calls for data trusts and unions as institutional forms of collective governance (delacroix2019bottom), we frame our approach as a form of algorithmic collective action in which contributors act collectively through a Data Union (DU) that sets group data disclosure policies, shifting how value and costs are distributed. While recent work shows that differentially-private mechanisms  dwork2006differential can diminish the leverage of collective action in gradient-based training (solanki2025crowding), we instead use DP positively as a tool for disclosure and value control.
@@
Figure 1: Illustration of the Information Disclosure Game. A Data Union (DU) holds a private dataset and releases information to a Data Consumer (DC) by adding Laplacian noise to data points under $\epsilon$-differential privacy. The DC incrementally acquires these noisy version of the data points, denoises them using an average and train a model to reach a utility target. In this work, we focus on non-parametric $k$-Nearest Neighbors ($k$-NN) and SBERT embeddings reimers-2019-sentence-bert.
We formalize this interaction as an Information Disclosure Game (IDG), a form of Stackelberg game simaan1973stackelberg; fudenberg1991game in which the DU acts as the leader by choosing disclosure rules, and the DC acts as the follower by optimizing acquisitions given those rules (Figure 1). This framework allows for the explicit identification of the acquisition cost that valuation imposes on the DC and how collective disclosure policies shape their distribution across members. We instantiate the model through a $k$-Nearest Neighbors ($k$-NN), in which the relationship between individual points and predictive utility is direct. Simulating the DC strategies guided by data Shapley values and by multi-armed bandit exploration, we evaluated it on a Yelp review helpfulness prediction task using text embeddings.
@@
First in Section 2, we review the related work on data valuation and differential privacy before detailing in Section 3 the IDG framework and the objectives of DU and DC. Afterwards, in Section 4, we present our empirical evaluation followed by a discussion and conclusion.
@@ 2 Background
In cooperative game theory, an imputation is a way to distribute the total value of a game among its players osborne1994course. In the context of data valuation, each data point in a dataset $D$ is treated as a player and a value function $v:2^{D}\to\mathbb{R}$ is used to the utility (e.g., accuracy or loss reduction) of any subset $D^{\prime}\subseteq D$. The objective is to find an imputation that allocates fairly the total value $v(D)$ among individual data points based on their contributions. Several approaches have been explored for data valuation—each with different fairness, computational and interpretability trade-offs—including Data Shapley ghorbani2019data; jia2019towards, Data Banzhaf wangDataBanzhafRobust2023, Beta Shapley kwonBetaShapleyUnified2022 and the Core yanIfYouShapley2021a.
@@
in which $D^{\prime}$ is a subset of $D\setminus\{z_{i}\}$ sampled uniformly at random and $v(D^{\prime})$ is the value function evaluated on $D^{\prime}$. The Shapley value thus quantifies the expected incremental contribution of each data point across all possible subsets. The uniqueness of the Shapley value can be derived from satisfying four foundational axioms: Efficiency (i.e., the total value is distributed among all players), Symmetry (i.e., identical contributions are rewarded equally), Dummy (i.e., players that contribute nothing receive zero value) and Linearity (i.e., the Shapley values from two games can be combined linearly). One application is data pricing in marketplaces jiaEfficientDataValuation2019; peiSurveyDataPricing2022; tianPrivateDataValuation2022; xiaEquitableDataValuation2023, in which Shapley-based frameworks assign higher prices to points that contribute across many subsets. Shapley values can also inform the data selection task, by identifying the examples that are the most useful for training smaller high-quality sets. Their effectiveness depends on the structure of the utility function and faces the challenge of a high computational cost, with approximations such as G-Shapley ghorbani2019data; yoon2020data proposed to improve scalability. However, for $k$-NN classifiers, an exact Shapley formulation exists jia2019efficient with complexity $O(N\log N)$, making valuation tractable in this setting.
@@
Differential Privacy (DP) dwork2006differential provides a mathematically rigorous framework for limiting how much information about an individual can be inferred from a released output. In this work, we focus on the local setting, in which each data point is randomized before it is disclosed to any external party cormode_privacy_2018; wang2019collecting. Formally, let $z$ be an individual data point and let $R$ be a randomized algorithm acting on single records. We say that $R$ satisfies $\epsilon$-local differential privacy if, for any pair of possible inputs $z,z^{\prime}$ and any measurable set of outputs $O$,
$\Pr[R(z)\in O]\leq e^{\epsilon}\Pr[R(z^{\prime})\in O].$
Classically, local DP is implemented in a decentralized manner, by having each individual perturbing his own data before sending it to any curator. Our setting is a bit different as the Data Union acts as a trusted curator and centrally applies a LDP-mechanism at the record level by adding noise independently to each member’s data point before releasing it to the Data Consumer. Beyond protection, DP is also related to the incentives individuals face when deciding whether to share their data truthfully. From a mechanism design perspective, the Revelation Principle myerson1983mechanism states that outcomes achievable through indirect strategies can also be implemented by truthful revelation. DP makes such mechanisms approximately truthful mcsherry2007mechanism, an $\varepsilon$-DP mechanism bounds each agent’s influence on the outcome, so the gain from lying is at most $O(\varepsilon)$.
@@ 3 Information Disclosure Game
We model the interaction between two agents: the DU and the DC. More precisely, the DU manages pooled data with the mission of ensuring fair valuation while accounting for concerns such as privacy. Meanwhile, the DC seeks to optimize utility—often by acquiring data points at minimal cost. These interactions are formalized as a two-phase Stackelberg game, in which the DU acts as a leader by setting how to disclose information and at what cost, and the DC, as a follower, responds through strategic data acquisition.
@@
This approach provides a straightforward pricing mechanism but has significant limitations. Indeed, as the disclosure of a data point is binary—either withheld or complete—many contributors receive no compensation. Incentivizing broader acquisition requires lowering the price of high-value points, which penalizes those who contribute the most. Moreover, releasing data at fixed prices locks in value at a single point in time and weakens privacy guarantees, limiting the DU’s ability to balance inclusiveness and protection.
@@
To address these challenges, we depart from per-point pricing and instead model an iterative disclosure process under DP. At each round $t$, the DC selects a subset $S_{t}$ of data points to query. Each query consumes a fixed privacy budget $\epsilon$ and returns a noisy version of the data. The DU sets $(\epsilon,T_{\max})$, the per-query budget and the per-point query cap, which together determine a maximum per-point spend $B_{\max}=T_{\max}\epsilon$. The DC’s cumulative spend is $B=\epsilon\sum_{t=1}^{T}|S_{t}|,\quad T\leq T_{\max}$. The DC’s objective is to minimize its cumulative budget spent:
@@
$\max_{(\epsilon,T_{\max})}\;\min_{\{S_{t}\}}\epsilon\sum_{t=1}^{T}|S_{t}|,$ (3)
with the goal of spreading being spent more evenly across members while managing privacy–utility trade-offs. Compared to the pricing game in Equation 1, partial disclosure replaces per-point ownership costs with iterative access charges. This formulation makes the hidden acquisition cost of valuation explicit in the DC’s budget and allows the DU to shape acquisition behaviors without penalizing the most valuable contributors.
@@ 4 Experiments on Review Helpfulness Prediction
Review helpfulness prediction provides a natural testbed for data valuation as platforms such as Yelp collect reviews not only for sentiment analysis but also to assess their quality. Indeed, unhelpful reviews may ultimately be discarded, yet they are still necessary to identify and reward helpful ones. This creates a setting in which individual contributions to predictive performance are both explicit and uneven. We use the Yelp dataset asghar2016yelp following the setup of bilal2023effectiveness, who report a $k$-NN accuracy of 59.6%. Using pretrained text embeddings and $k$-NN, our best configuration achieves 66.0% test accuracy and 65.2% F1-score, bringing non-parametric performance closer to fine-tuned transformer models while preserving point-level interpretability needed for valuation. As a sanity check, we replicated the acquisition experiment of jia2019towards and verified that Shapley-based selection outperforms random sampling. The corresponding curve, along with additional details on the experimental setup, is reported in Appendix D.
@@
in which $\Delta_{j}=b_{j}-a_{j}$ denotes the global sensitivity of feature $j$ after bounding and $\epsilon_{j}$ is the allocated budget for that feature. This design is similar in spirit to a per-feature local DP mechanism cormode_privacy_2018; wang2019collecting, in that each coordinate is privatized independently, but here it is applied in a centralized setting by the DU.
We assume the sensitivity interval $a_{j},b_{j}]$ is fixed in advance and does not vary significantly across releases, so the noise scale is determined by $\Delta_{j}/\epsilon_{j}$. On the DC side, each noisy version of a point is averaged to form a denoised estimate. After observing $t$ noisy versions of point $i$, its center is computed incrementally as $\hat{x}_{i}^{(t)}=\frac{1}{t}\sum_{k=1}^{t}\tilde{x}_{i}^{(k)}$. The averaging strategy steadily improves fidelity as noisy samples accumulate, with error decreasing and correlation increasing non-linearly across iterations (see Appendix B for details).
@@ 4.1 Data Selection Using Random and Shapley-based Strategies
In this baseline, the DC commits to a random subset of data points for all iterations, with the results displayed in Figure 3 being averaged over 10 random permutations. As shown in Figure 3, this approach fails to reach the target utility within 100 iterations unless nearly 100% of the dataset is acquired, which confirms that a purely random acquisition strategy from the DU is not viable under budget constraints.
@@ 4.2 Data Selection using $n$-Armed Bandits
The utility of a queried point is defined as the fraction of neighbors in $k$-NN that share its label. See Algorithm 1 in appendices for implementation details.
@@
Identifying an equilibrium can be seen as a hyperparameter search, since the interaction between the DU’s release policy and the DC’s acquisition strategy does not admit a closed-form solution. To realize this, we conducted a grid search over two key parameters: the maximum budget per point (set by the DU) and the exploration coefficient $c$ (in UCB). These jointly determine budget usage and the diversity of selected points. Results show that the bandit reliably reaches the utility threshold except when the per-point budget is too small or exploration is disabled; below a budget of 20, success is rare and often due to lucky point selection. Gini analysis of budget allocations confirms that exploration promotes more balanced spending across points and increases the likelihood of success. The correlation between Q values and Shapley values grows with the budget but remains moderate overall (Appendix E).Finally, evaluations based on denoised centers yielded test accuracy comparable to or exceeding that obtained from the original data. For instance, a configuration with a per-point budget of 50 reached 69.9% surpassing the 66% benchmark.
@@ 5 Discussion & Conclusion
We proposed a game-theoretic framework for data valuation that integrates privacy and inclusiveness concerns through an information disclosure game between a Data Union (DU) and a Data Consumer (DC). Rather than assuming direct access to data, our approach models acquisition as an iterative process using a differentially-private mechanism to enforce a cost on data valuation. On Yelp review helpfulness prediction, Shapley-based strategies remained effective under noise but required a minimal budget—about 10 bootstrap iterations—to reliably identify high-value points. This demonstrates that Shapley valuation entails an inherent cost of exploration in our setting. Similarly, a multi-armed bandit strategy reached the utility target with comparable budgets, reinforcing that valuation itself imposes exploration costs. Gini analysis further showed that as budgets grow, the DC allocates resources more evenly across contributors, increasing inclusiveness.
While our experiments focused on $k$-NN for computation efficiency, an important direction for future work is extending the framework to differentiable models. Gradient-based methods such as G-Shapley ghorbani2019data already show that Shapley values can be approximated from changes in gradients, suggesting ways to scale valuation beyond non-parametric models. In parallel, approaches like DP-SGD abadi2016deep demonstrate that DP can be enforced by adding noise to gradients rather than directly to points. Bridging these lines of work would provide a better privacy-utility trade-off for both the DU and DC. Nonetheless, our results suggest that in a DU setting, reaching target utility requires a minimum level of budget spread. Thus, our framework implies that a minimum dividend should be guaranteed to all members, regardless of their individual Shapley value. In the context of reviews, this means that every contributor would receive at least some share of value, even if their individual review is ultimately deemed unhelpful, aligning the incentives of all members with the collective outcome. Otherwise, excluding some contributors increases the incentive to form data unions that adopt adversarial disclosure strategies (e.g., prioritizing low-value reviews or injecting excess noise), which may in turn hinder the data consumer’s ability to reach its utility target.